OpenStudy (anonymous):

Please help I don't know what I did wrong Solve for x: |x + 2| + 16 = 14

OpenStudy (cwrw238):

|x + 2| = 14 - 16 = -2 did you get this?

OpenStudy (anonymous):

i got that x = -4 and 0

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

thats incorrect but did you get what i got in my post?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution those are the answer choices

OpenStudy (cwrw238):

do you know what the vertical lines around the x + 2 mean?

OpenStudy (anonymous):

yes absolute value

OpenStudy (cwrw238):

right - and the absolute value is always POSITIVE for example |- 2| = 2 and we got |x + 2| = -2 which cannot be true

OpenStudy (cwrw238):

the absolute value cannot be negative

OpenStudy (anonymous):

so its no solution

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (cwrw238):

yw